OpenStudy (dannyrod2000):

sqaure roots of 100. Check all that apply.

OpenStudy (dannyrod2000):

http://prntscr.com/4bvkeg

OpenStudy (dannyrod2000):

C,D,and F?

OpenStudy (dannyrod2000):

@e.mccormick

OpenStudy (dannyrod2000):

@Hero

OpenStudy (anonymous):

You got it right! C, D, and F

OpenStudy (dannyrod2000):

so i right?

OpenStudy (dannyrod2000):

got it right! :)

OpenStudy (anonymous):

10*10=100 -10*-10=100, because a negative times a negative is a positive |10|*|10|=100, because the absolute value of 10 is 10.

OpenStudy (dannyrod2000):

xD i knew it :)

OpenStudy (anonymous):

Did that help explain?

OpenStudy (dannyrod2000):

yea thats what i knew

OpenStudy (anonymous):

Awesome! Good luck on the rest, and message me if you need help!

OpenStudy (dannyrod2000):

k thanks

OpenStudy (dannyrod2000):

can i tag you?